OpenStudy (anonymous):

Find the exact value of sin2u under the given conditions. tan(u)=2/5 abd 180 degrees < u < 270 degrees

OpenStudy (raden):



OpenStudy (raden):

first, u have to find the length of base that triangle by using pythagorean theorem

OpenStudy (anonymous):

sin2u=2sinucosu=2sinucosu/((cosu)^2+(sinu)^2)=2/(tanu+1/tanu)=20/29 u get (20,21,29)